OpenStudy (anonymous):

Write the quadratic variation equation for the relationship: y varies directly with x2 and y = 36 when x = 3.

OpenStudy (oaktree):

If y=k*x^2, then 36=k*(3)^2, 36=9k, and k=4. So the equation is y=4x^2.